# Why Isn’t Relational Learning Taking Over the World?

### v2

@@ Introduction
[…]
@@
  • Does the system know what exists? Most databases assume that they know what exists; things that exist are given identifiers, and what doesn’t exist is ignored. In many real-world cases, we don’t know what exists. What there a person in the house last night? Who threw the rock that broke the window? (It might have not been a rock or maybe no one threw it.)
@@
One large public knowledge graph is Wikidata (Vrandečić and Krötzsch 2014). Figure 2 (a) shows some tables representing the football (soccer) player Christine Sinclair. She is the player (man or woman) with the most goals in international play. Figure 2 (a) shows the corresponding Wikidata knowledge graph. In Wikidata, Christine Sinclair is represented by the identifier Q262802. The country Canada is represented by the identifier Q16.
@@
Including only most frequent entities means that there are no reified entities in the resulting triples. In Wikidata, around 99% of the entities appear as the subject of fewer than 10 triples (see Appendix A). Training and evaluating on datasets where all such entities are removed will not lead to better predictors on more realistic datasets.
Big data often implies more entities, but less data, on average, about each entity. As entities are added, they start off as stubs with few tuples, resulting in a long tail of entities about which not much is known. As more detail is added about an entity, there are more associated reified entities, about which there are inherently few tuples. See Appendix A.
@@
[…]
@@
[…]
@@
[…]
@@
When predicting entities, rather than assuming that the correct answer is one of the known entities, there are three possible types of answers: one or more of the known entities, an entity or entities not represented, or that there is no entity. Consider asking for birth mother. For most people in a knowledge base the answer is “someone not represented”. Asking for a child of a person, a person may have (perhaps multiple) children represented, children not represented (or both) or have no children. Similarly, asking for an event that caused something, it might be a known event, an unknown event or there might no event. An unknown event may need to be given an identifier to give it properties. A prediction should not be judged on the actual identier assigned.
We need to be clear about whether we want to predict general knowledge or learn about the entities in the training set. Unless the test sets are carefully designed, there will tend to be a bias towards learn about just one of these. Many models such as probabilistic logic programs (Poole 1993; De Raedt, Kimmig, and Toivonen 2007) and convolutional graph neural networks (Schlichtkrull et al. 2018) learn what looks like general knowledge that is then be applied to the training set to predict properties and relations of the entities in training set. Tests sets from the same dataset do not test whether these models generalize to new populations or related domains. To have statistics, we need to combine the data in at least one dimension; e.g., we could combine temporal data to learn about road congestion at a particular intersection, but could also generalize over the type of intersection. Map builders build models that generalize over only some of the dimensions, but I haven’t seen it for relational learning in general.
@@
To realize the full potential of relational learning, we need to combine the information from all datasets, and build predictive models from these. Combining multiple heterogenous datasets is the challenge of the semantic web (Berners-Lee, Hendler, and Lassila 2001). Building predictive models from heterogenous data might seem impossible, but we know how to; it’s called science. To combine all recorded experimental and observational data to make better predictions, we need at least ontologies [^6] Scientific ontologies include Open Biological and Biomedical Ontology http://obofoundry.org, WHO Family of International Classifications https://www.who.int/standards/classifications, SMOMED-CT healthcare ontology https://www.snomed.org (to allow semantic interoperability) and provenance information (Gil et al. 2016; Sikos, Seneviratne, and McGuinness 2021), including what was manipulated, what was controlled for, what was observed and when, how was what to record decided (modelling the missing information), who manipulated the data, etc. Science involved building hypotheses, typically from multiple datasets, that persist, ready to be revised or rejected as new evidence comes in. Hypotheses can be about anything; from the effects of global warming, to what foods are edible, to the traffic in Vancouver. There needs to be multiple hypotheses for a dataset, including the null hypothesis (the data is noise) and hypotheses that some data is erroneous or entirely fictitious (to counter misinformation). Multiple hypotheses may need to be combined to make a prediction for an actual case (Poole 2012). With an interconnected world, we don’t need to make conclusions from individual experiments, but can learn from everyones experiences, which should enable more accurate hypotheses.
@@ Appendix A Appendix A: Proportion of Wikidata entities with few tuples
The aim of this appendix is to estimate the proportion of entities in Wikidata that are the subject of few tuples. There is very long tail of entities about which there are few tuples, and presumably are stubs that will be expanded in the future, so it is unfair to count those. Instead we approximate the number of reified entities per top-level entity. I could not download all of Wikidata, so as a sample I chose the entities of the Wikipedia featured articles of March 2025 (https://en.wikipedia.org/wiki/Wikipedia:Today%27s˙featured˙article/March˙2025) that had Wikidata pages. The entities were diverse including Hughie Ferguson, a Scottish footballer in 1910s and 20s, Pulgasari, a North Korean film from 1985, the contemporary ballet Flight Pattern, Steele’s Greenville expedition in the American Civil War, the sun, and a fictional character in the medical drama series House.
@@
For each entity (including reified entities), Wikidata allows the user to download the triples about that entity. For reified entities, the triples downloaded are the triples of the subject of the reification. We define the local entities for entity $e$ to be those that appear as a subject in a triple for the page for $e$ and for which Wikidata downloads the page for $e$. For each of the selected entities, we counted the number of local entities that are the subject of fewer than 10 triples. See Figure 4. The number ranged from 5 for History of infant schools in Great Britain and 6 for Your Girl (a 2005 song by Mariah Carey), to 836 for the Sun and 506 for the composer and conductor Pierre Boulez. The lowest of these are just stubs; there are only 11 and 8 tuples respectively with the entity as the subject of a tuple (the 11 included the name and labels in both English and Portuguese). The average number of local entities was 129 (or 108 if the top and bottom outliers were removed). If, for each entity, there are 108 entities that are the subject of fewer than 10 triples, then these entities are 99.08% of the entities.
Wikidata ID # tuples #<10\#<10# < 10 name
Q107132670 67 34 Beverly White
Q3142353 110 50 Hughie Ferguson
Q931229 87 38 USS Congress
Q11813 1055 471 James Madison
Q3830799 157 60 Leroy Chollet
Q6529597 110 49 Les Holden
Q133116210 11 5 History of infant schools in Great Britain
Q4767002 208 89 Anna Filosofova
Q485172 679 283 Ann Arbor
Q16267674 397 277 Hotline Miami 2: Wrong Number
Q1349580 198 92 Matthew Brettingham
Q632511 302 103 Seattle Sounders FC
Q1137583 237 118 Pulgasari
Q3518834 58 20 Territorial Force
Q98076985 29 21 Flight Pattern
Q822900 211 36 geography of Ireland
Q192920 463 160 Edward the Martyr
Q116245633 19 17 Steele’s Greenville expedition
Q525 2588 835 Sun
Q122330832 66 47 All-American Bitch
Q130337417 32 17 The True Record
Q3856283 69 29 Michael Tritter
Q556464 268 156 Gertie the Dinosaur
Q5461994 177 110 Flotilla
Q156193 845 506 Pierre Boulez
Q545151 140 51 The Spy Who Loved Me
Q830590 62 32 Interstate 182
Q3142456 75 41 Hurricane Cindy
Q131151772 9 6 Your Girl
Average 301 129
Figure 4: For each entity, the number of tuples with the entity as subject, the number of local entities that are the subject of fewer than 10 tuples, and the English name of the entity
Thus for each top-level entity there about 100 other entities with very few tuples, and is of the right order-of-magnitude. There are also top-level entities with few triples. Thus, 99% seems like a reasonable estimate.
@@
As the number of triples increases for an entity, so does the number of entities with few triples. Thus, as a knowledge graph gets bigger and more filled out, the proportion of entities with few triples increases. This goes against the idea that we just need more data.
@@ Appendix B Appendix B: An MRR agent
Companies that present lists of results, such as Google, will know the distribution of the number of items considered. It would be better to use this distribution for evaluation than MRR, but it probably depends on what is being presented (and whether predictions can be tested to determine if that are correct or what is wanted).

### v3

@@ Introduction
where Next is the next token, Con, the context, is the prompt and the previously generated text, and acc is the proposition that the text is acceptable from human feedback. $P(\mbox{\emph{Next}}\mid\mbox{\emph{Context}})$ is trained from enormous datasets. $P(\mbox{\emph{acc}}\mid\mbox{\emph{Next}},\mbox{\emph{Context}})$, reinforcement learning by human feedback (RLHF) [^1] To see how this is RLHF, $Acc$ is a reward of 1 for acceptable and 0 for unacceptable. Expected utility is equivalent to probability of Acc=1. Humans evaluate complete text. $P(\mbox{\emph{Acc}=1}\mid\mbox{\emph{Next}},\mbox{\emph{Context}})$ is the $Q$-function of reinforcement learning, where Next corresponds to the action and Context is the state. (Knox and Stone 2009; Ziegler et al. 2020), is learned from humans evaluating generated text.
@@
  • Does the system know what exists? Most databases assume that they know what exists; things that exist are given identifiers, and what doesn’t exist is ignored. In many real-world cases, we don’t know what exists. Was there a person in the house last night? If so, who? Who threw the rock that broke the window? (It might have not been a rock or maybe no one threw it.)
@@
One large public knowledge graph is Wikidata (Vrandečić and Krötzsch 2014). Figure 2 (a) shows some tables representing the football (soccer) player Christine Sinclair. She is the player (man or woman) with the most goals in international play. Figure 2 (b) shows the corresponding Wikidata knowledge graph. In Wikidata, Christine Sinclair is represented by the identifier Q262802. The country Canada is represented by the identifier Q16.
@@
Including only most frequent entities means that there are no reified entities in the resulting triples. In Wikidata, over 98% of the entities appear as the subject of fewer than 10 triples (see Appendix A). Training and evaluating on datasets where all such entities are removed will not lead to better predictors on more realistic datasets.
Big data often implies more entities, but less data, on average, about each entity. As entities are added, they start off as stubs with few tuples, resulting in a long tail of entities about which not much is known. As more detail is added about an entity, there are more associated reified entities, about which there are inherently few tuples (see Appendix A).
@@
The evaluation of a knowledge graphs depends on what is asked. To predict whether a random triple is in Wikidata, answering “no” is over 99.99999999995% accurate [^4] Wikidata has about 1.65 billion triples on 117 million items and 290,000 properties (https://grafana.wikimedia.org/d/000000175/wikidata-datamodel-statements). With no domain or range constraints in Wikidata, any entity can be the subject or object of any property. (Let’s ignore strings and other datatypes as values, as there are too many possible values). There are thus $(1.17e8)^{2}*2.9e5\approx 4e21$ possible triples, of which 1.6e9/4e21=4e131.6e9/4e21=4e{-}13 are in Wikidata. . Accuracy of random triples is not a reasonable measure. Most random triples are nonsense, which no one would ask and could be answered just by knowing the types of the entities.
@@
The most common way to evaluate a knowledge graph predictor is to use ranking: given test triple $(s,v,o)$, ask the predictor to rank the entities related to a given subject $s$ and verb $v$, i.e., to predict which triples of the form $(s,v,?)$ are true (and similarly to rank the subjects for a given verb and object). Given a total ordering of predictions $o_{1},o_{2},o_{3},\dots$, the rank is the $i$ such that $o=o_{i}$. In filtered prediction (Bordes et al. 2013b), the $o_{i}$ that are known to be true (because the corresponding triple is in the training or test set) are removed from the ordering.
The mean rank is one possible measure but is dominated by a few very large ranks, and is not used much any more. A common measure is “hit-at-kk” for some integer $k$ (such as 1, 3 or 10), the proportion of test cases where the rank of the correct answer is $k$ or less. The $k$ in hit-at-$k$ seems like an arbitrary parameter. A method that is common and seems to not have arbitrary parameters is the mean reciprocal rank (MRR) (Radev et al. 2002). This weights the $k$th rank as $1/k$. It can be interpreted as an agent who always considers the first item, considers the second item half the time, the 3rd item 1/3 of the time on so on. However this interpretation is problematic (see Appendix B).
@@
  • It loses sight of the downstream task. Considering top-$k$ is reasonable for cases where the results can be checked, such as in movie recommendation, web search or suggesting to a doctor the $k$ most likely diagnoses so they can be reminded of something they may have overlooked. It seems less useful when just presenting $k$ possible answers for which there is no easy way to check which, if any, are correct. State-of-the-art methods that combines existing models (Li et al. 2024; Jeon, Lim, and Suk Choi 2024) have hit-at-10 rates on FB15k-237 of 55.8%. and 52.58% respectively, which does not seem very useful for any real-world task where correctness cannot be easily verified.
@@
When predicting entities, rather than assuming that the correct answer is one of the known entities, there are three possible types of answers: one or more of the known entities, an entity or entities not represented, or that there is no entity. Consider asking for birth mother. For most people in a knowledge base the answer is “someone not represented”. Asking for a child of a person, a person may have (perhaps multiple) children represented, children not represented (or both) or have no children. Similarly, asking for an event that caused something, it might be a known event, an unknown event or there might no event. An unknown event may need to be given an identifier to give it properties. A prediction should not be judged on the actual identifier assigned.
We need to be clear about whether we want to predict general knowledge or learn about the entities in the training set. Unless the test sets are carefully designed, there will tend to be a bias towards learn about just one of these. Many models such as probabilistic logic programs (Poole 1993; De Raedt, Kimmig, and Toivonen 2007) and convolutional graph neural networks (Schlichtkrull et al. 2018) learn what looks like general knowledge that is then be applied to the training set to predict properties and relations of the entities in training set. Tests sets from the same dataset do not test whether these models generalize to new populations or related domains.
To have statistics, we need to combine the data in at least one dimension. E.g., we could ask for the average congestion at a particular intersection at a particular time, or the average congestion at intersections on a particular date and time, but not at at a particular intersection at a particular date and time, as there is at most a single data point. There could be latent features (embeddings) for intersections and dates, but some intersections and dates are peculiar and it might be better to learn about particular entities and not just general knowledge.
@@
To realize the full potential of relational learning, we need to combine the information from all datasets, and build predictive models from these. Combining multiple heterogenous datasets is the challenge of the semantic web (Berners-Lee, Hendler, and Lassila 2001). Building predictive models from heterogenous data might seem impossible, but we know how to; it’s called science. To combine all recorded experimental and observational data to make better predictions, we need at least ontologies [^6] Scientific ontologies include Open Biological and Biomedical Ontology http://obofoundry.org, WHO Family of International Classifications https://www.who.int/standards/classifications, SMOMED-CT healthcare ontology https://www.snomed.org (to allow semantic interoperability) and provenance information (Gil et al. 2016; Sikos, Seneviratne, and McGuinness 2021), including what was manipulated, what was controlled for, what was observed and when, how was what to record decided (modelling the missing information), who manipulated the data, etc. Science involves building hypotheses, typically from multiple datasets, that persist, ready to be revised or rejected as new evidence comes in. Hypotheses can be about anything; from the effects of global warming, to what foods are edible, to the traffic in Vancouver. There needs to be multiple hypotheses for a dataset, including the null hypothesis (the data is noise) and hypotheses that some data is erroneous or entirely fictitious (to counter misinformation). Multiple hypotheses may need to be combined to make a prediction for an actual case (Poole 2012). With an interconnected world, we don’t need to make conclusions from individual experiments, but can learn from everyones experiences, which should enable more accurate hypotheses.
@@ Appendix A Appendix A: Proportion of Wikidata entities with few tuples
The aim of this appendix is to estimate the proportion of entities in Wikidata that are the subject of few tuples. There is very long tail of entities about which there are few tuples, and presumably are stubs that will be expanded in the future, so it may be unfair to count those. Instead, we approximate the number of reified entities per top-level entity. I could not download all of Wikidata, so as a sample I selected the entities of the Wikipedia featured articles of March 2025 (https://en.wikipedia.org/wiki/Wikipedia:Today%27s˙featured˙article/March˙2025) that had Wikidata pages. The entities were diverse including Hughie Ferguson, a Scottish footballer in 1910s and 20s, Pulgasari, a North Korean film from 1985, the contemporary ballet Flight Pattern, Steele’s Greenville expedition in the American Civil War, the sun, and a fictional character in the medical drama series House.
@@
For each entity (including reified entities), Wikidata allows users to download triples about that entity. For reified entities, the triples downloaded are the triples of the subject of the reification. We define the local entities for entity $e$ to be those that appear as a subject in a triple for the page for $e$ and for which Wikidata downloads the page for $e$. For each of the selected entities, we counted the number of local entities that are the subject of fewer than 10 triples and the number of entities that are the subject of 10 or more triples. See Figure 4. The number ranged from 5 for History of infant schools in Great Britain and 6 for Your Girl (a 2005 song by Mariah Carey), to 841 for the Sun and 520 for the composer and conductor Pierre Boulez. The entities with the lowest number of tuples are just stubs; note that for Q133116210, the 11 tuples included the name and labels in both English and Portuguese. The average number of local entities was 134 (or 108 if the top and bottom outliers were removed).
Wikidata ID # tuples <10<10 $\geq 10$ name
Q107132670 70 34 1 Beverly White
Q3142353 125 50 5 Hughie Ferguson
Q931229 87 38 1 USS Congress
Q11813 1074 490 10 James Madison
Q3830799 161 63 1 Leroy Chollet
Q6529597 116 49 1 Les Holden
Q133116210 11 5 1 History of infant schools in Great Britain
Q4767002 210 89 1 Anna Filosofova
Q485172 684 285 6 Ann Arbor
Q16267674 426 282 3 Hotline Miami 2: Wrong Number
Q1349580 201 95 2 Matthew Brettingham
Q632511 325 113 4 Seattle Sounders FC
Q1137583 242 121 1 Pulgasari
Q3518834 58 20 1 Territorial Force
Q98076985 29 21 1 Flight Pattern
Q822900 214 36 1 geography of Ireland
Q192920 475 172 2 Edward the Martyr
Q116245633 19 17 1 Steele’s Greenville expedition
Q525 2605 841 2 Sun
Q122330832 66 47 2 All-American Bitch
Q130337417 32 17 1 The True Record
Q3856283 70 30 1 Michael Tritter
Q556464 270 158 1 Gertie the Dinosaur
Q5461994 201 112 1 Flotilla
Q156193 859 520 5 Pierre Boulez
Q545151 141 51 2 The Spy Who Loved Me
Q830590 62 32 1 Interstate 182
Q3142456 78 41 1 Hurricane Cindy
Q131151772 9 6 0 Your Girl
Average 308 132 2
Figure 4: For each top-level entity: the Wikidata ID, the number of tuples with the entity as subject, the number of local entities that are the subject of fewer than 10 tuples, the number with 10 or more tuples, and the English name of the entity. These numbers were from 2025-08-19; they change over time as the knowledge base evolves. The Python code to generate the contents of this table is available from the author.
The entities that are the subject of 10 or more tuples (other than the top-level entity) are refied entities from relations with a large number of arguments; the maximum of which is 17. About 1.5% of the entities are the subject of 10 or more triples. Fewer than 0.75% of the entities are the subject of 18 or more triples.
@@
As the number of triples increases for an entity, so does the number of entities with few triples. Thus, as a knowledge graph gets bigger and more filled out, the proportion of entities with few triples increases. This goes against the idea that we just need more data. There are also top-level entities with few triples.
@@ Appendix B Appendix B: An MRR agent
Companies that present lists of results and get feedback about which item(s) was selected or how many items were presented, such as Google, will have data on the distribution of the number of items considered. It would be better to use this distribution for evaluation than MRR, but it probably depends on what is being presented (and whether predictions can be tested to determine if that are correct or what is wanted).